# PoseAnimate: Zero-shot high fidelity pose controllable character animation

### v1

@@ 3.3 Dual Consistency Attention Module
Although the pose-aware control module accurately captures and injects body poses, it may unintentionally alter the identity of the character and the background details due to the introduction of different pose signals, as demonstrated by the example $\widetilde{Z}_{x_{i},0}$ in Fig. 3, which is undesirable. Since self-attention layers in the U-Net ronneberger2015u play a crucial role in controlling appearance, shape, and fine-grained details, existing attention fusion paradigms commonly employ cross-frame attention mechanism ni2022expanding , to facilitate spatial information interaction across frames:
@@
where $Q^{i}$ is the query feature of frame $x_{i}$, and $K^{j},V^{j}$ correspond to the key feature and value feature of frame $x_{j}$. As pose $p_{1}$ is identical to the original pose $p_{s}$, the reconstruction of frame $x_{0}$ remains undisturbed, allowing for a perfect restoration of the source image $I_{s}$. Hence, we can compute the cross-frame attention between each subsequent frame $\{x_{i}\}_{i=1}^{N}$ with the frame $x_{0}$ to ensure the preservation of identity and intricate details. However, solely involving frame $x_{0}$ in the attention fusion would bias the generated actions towards the original action, resulting in ghosting artifacts and flickering. Consequently, we develop the Dual Consistency Attention Module(DCAM) by replacing self-attention layers with our dual consistency attention(DC Attention) to address the issue of appearance inconsistency and improve temporal consistency. The DC Attention mechanism operates for each subsequent frame $x_{i}$ as follows:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}\in(0,1)$ are hyper-parameters, and $\lambda_{1}+\lambda_{2}+\lambda_{3}=1$. $\text{CFA}_{i,j}$ refers to cross-frame attention between frames $x_{i}$ and $x_{j}$.They jointly control the participation of the initial frame $x_{0}$, the current frame $x_{i}$ and the preceding frame $x_{i-1}$ in the DC Attention calculation. In the experiment, we set $\lambda_{1}=0.7$ and $\lambda_{2}=\lambda_{3}=0.15$ to enable the frame $x_{0}$ to be more involved in the spatial correlation control of the current frame for the sake of better appearance preservation. Apart from this, retaining a relatively small portion of feature interaction for the current frame and the preceding frame simultaneously is promised to enhance motion stability and improve temporal coherence of the generated animation.
Furthermore, it is vital to note that we do not replace all the U-Net transformer blocks with DCAM. We find that incorporating the DC Attention only in the upsampling blocks of the U-Net architecture while leaving the remaining unchanged allows us to maintain consistency with the identity and background details of the source, without compromising the current frame’s pose and layout.
@@ 3.4 Mask-Guided Decoupling Module
Directly utilizing the entire image features for attention fusion can lead to substantial loss of fine-grained details. To address this problem, we propose the mask-guided decoupling module, which decouples the character and background and enables individual inter-frame interaction to further refine spatial feature perception.
For the source image $I_{s}$, we obtain a precise body mask $M_{s}$ (i.e. $M_{x_{0}}$) that separates the character from the background by an off-the-shelf segmentation model liu2023lightweight . The target pose prior is insufficient to derive body mask for each generated frame of the character. Considering the strong semantic alignment capability of cross attention layers mentioned in Prompt-to-prompt hertz2022prompt , we extract the corresponding body mask $M_{x_{i}}$ for each frame from the cross attention maps. With $M_{s}$ and $M_{x_{i}}$, only attentions of character and background within corresponding region are calculated, according to the mask-guided decoupling module as follows:
$\text{K}^{c}_{j}=M_{x_{j}}$ $\odot\text{K}_{j},\text{K}^{b}_{j}=(1-M_{x_{j}})\odot\text{K}_{j}$ (8)
$\text{V}^{c}_{j}=M_{x_{j}}$ $\odot\text{V}_{j},\text{V}^{b}_{j}=(1-M_{x_{j}})\odot\text{V}_{j}$
@@
$\text{DCA}^{c}_{i}$ $=\lambda_{1}*\text{CFA}^{c}_{i,0}+\lambda_{2}*\text{CFA}^{c}_{i,i-1}+\lambda_{3}*\text{CFA}^{c}_{i,i}$ (9)
$\text{DCA}^{b}_{i}$ $=\lambda_{1}*\text{CFA}^{b}_{i,0}+\lambda_{2}*\text{CFA}^{b}_{i,i-1}+\lambda_{3}*\text{CFA}^{b}_{i,i}$
@@ 4.1 Experiment Settings
We implement PoseAnimate based on the pre-trained weights of ControlNet zhang2023adding and Stable Diffusion rombach2022high v1.5. For each generated character animation, we generate $N=16$ frames with a unified $512\times 512$ resolution. All experiments are performed on a single NVIDIA A100 GPU.
@@ 4.2 Comparison Result
We compare our PoseAnimate with several state-of-the-art methods for character animation: MagicAnimate xu2023magicanimate and Disco wang2023disco . For MagicAnimate, both densepose guler2018densepose and openpose signals of the same motion are applied to evaluate performances. We leverage the official open source code of disco to test its effectiveness. Additionally, we construct a competitive character animation baseline by IP-Adapter ye2023ip with ControlNet zhang2023adding and spatio-temporal attention wu2023tune , which is termed as IP+CtrlN. It is worth noting that these methods are all training based, while ours does not require training.
@@
We set up two different levels of pose for experiments to fully demonstrate the superiority of our method. The visual comparison results are shown in Fig. 4, with the left side displaying simple actions and the right side complex actions. Although IP+CtrlN has good performance on identity preservation, it fails to maintain details and inter-frame consistency. Disco loses the character appearance completely, and severe frame jitter leads to ghosting shadows and visual collapse for complex actions. MagicAnimate performs better than the other two methods, but it still encounters inconsistencies in character appearance at a more fine-grinded level guided by Densepose. It is also unable to preserve background and character details accurately, e.g. vehicle textures and mask of firefighter and the boy in Fig. 4. MagicAnimate under OpenPose signal conditions has worse performances than that under DensePose. While our method exhibits the best performance on image fidelity to the source image, and effectively preserves complex fine-grained appearance details and temporal consistency.
@@
Figure 5: Ablation study. Source prompt: “An iron man on the road.”
@@
Table 1: Quantitative comparison between our PoseAnimate and other training-based state-of-the-art methods. The best average performance is in bold. ↑ indicates higher metric value and represents better performance and vice versa.
For quantitative analysis, we randomly sample 50 in-the-wild image-text pairs and 10 different disered pose sequences to conduct evaluations. We adopt four evaluation metrics: (1) LPIPS zhang2018unreasonable measures the fidelity between generated frames and source image. (2) CLIP-I ye2023ip represents the similarity of CLIP radford2021learning image embedding between generated frames and the source image. (3) Frame Consistency(FC) esser2023structure evaluates video continuity by computing the average CLIP cosine similarity of two consecutive frames. (4) Warping Error(WE) liu2023evalcrafter evaluates the temporal consistency of the generated animation through the Optical Flow algorithm teed2020raft .
Quantitative results are provided in Table. 1. Our method achieves the best scores on LPIPS and CLIP-I and greatly surpasses other comparison methods in terms of fidelity to the source image, demonstrating outstanding detail preservation capability. In addition, PoseAnimate outperforms the other two training-based methods in terms of inter-frame consistency. A good Warping Error score is also achieved, illustrating that our method is able to maintain good temporal coherence without additional training.
@@ 4.3 Ablation Study
We conduct ablation study to verify effectiveness of each component of our framework and present results in Fig. 5. The leftmost one in the first row is the source image, and the others are the target pose sequences. The following rows are generation results without certain components: (a) Pose-Aware Control Module (PACM) that effectively removes the interference of character original pose and maintains consistency of content unrelated to character; (b) Dual Consistency Attention Module (DCAM) that maintains image fidelity to the source image and improves temporal consistency; (c) Masked-Guided Decoupling Module (MGDM) that preserves image details; and (d) Pose Alignment Transition Algorithm (PATA) that tackles the issue of misalignments.
@@
Fig. 5(a) illustrates the significant interference of original pose on the generated actions. Due to the substantial difference between the posture of Iron Man’s legs in the source and in the target, there is a severe breakdown in the leg area of the generated frame, undermining the generation of a reasonable target action. Moreover, the character-irrelevant scenes also have noticeable distortion.
@@
From Fig. 5(b) we can find that it fails to maintain content consistency without Dual Consistency Attention Module. And the missing pole and Iron Man’s hand in the red box reveal inter-frame inconsistency, indicating that both spatial and temporal content cannot be effectively maintained.
@@
Compared with our results in Fig. 5(e), we can observe that small signs are missing without MGDM. It proves that Masked-Guided Decoupling Module can effectively enhance the fine-grained feature perception and image fidelity.
@@
Fig. 5(d) verifies the proposed Pose Alignment Transition Algorithm. The red circles in the first frame indicate the spatial content misalignment. When Iron Man in the original image does not match with the input pose position, an extra tree appears in the original position of Iron Man. And such misalignment can also leads to disappearance of background details, e.g., streetlights and distant signage.
@@ 5 Conclusion
This paper proposes a novel zero-shot approach PoseAnimate to tackle the task of character animation for the first time. PoseAnimate can generate temproal coherent and high-fidelity animations for arbitrary images under various pose sequences. Extensive experiment results demonstrate that PoseAnimate outperforms the state-of-the-art training based methods in terms of character consistency and detail fidelity.

### v3

@@ 3.3 Dual Consistency Attention Module
Although the pose-aware control module accurately captures and injects body poses, it may unintentionally alter the identity of the character and the background details due to the introduction of different pose signals, as demonstrated by the example $\widetilde{Z}_{x_{i},0}$ in Fig. 3, which is undesirable. Since self-attention layers in the U-Net Ronneberger et al. (2015) play a crucial role in controlling appearance, shape, and fine-grained details, existing attention fusion paradigms commonly employ cross-frame attention mechanism Ni et al. (2022), to facilitate spatial information interaction across frames:
@@
where $Q^{i}$ is the query feature of frame $x_{i}$, and $K^{j},V^{j}$ correspond to the key feature and value feature of frame $x_{j}$. As pose $p_{1}$ is identical to the source pose $p_{s}$, the reconstruction of frame $x_{0}$ remains undisturbed, allowing for a perfect restoration of the source image $I_{s}$. Hence, we can compute the cross-frame attention between each subsequent frame $\{x_{i}\}_{i=1}^{N}$ with the frame $x_{0}$ to ensure the preservation of identity and intricate details. However, solely involving frame $x_{0}$ in the attention fusion would bias the generated actions towards the original action, resulting in ghosting artifacts and flickering. Consequently, we develop the Dual Consistency Attention Module (DCAM) by replacing self-attention layers with our dual consistency attention (DC Attention) to address the issue of appearance inconsistency and improve temporal consistency. The DC Attention mechanism operates for each subsequent frame $x_{i}$ as follows:
@@
where $\text{CFA}_{i,j}$ refers to cross-frame attention between frames $x_{i}$ and $x_{j}$. $\lambda_{1},\lambda_{2},\lambda_{3}\in(0,1)$ are hyper-parameters, and $\lambda_{1}+\lambda_{2}+\lambda_{3}=1$. They jointly control the participation of the initial frame $x_{0}$, the current frame $x_{i}$ and the preceding frame $x_{i-1}$ in the DC Attention calculation. In the experiment, we set $\lambda_{1}=0.7$ and $\lambda_{2}=\lambda_{3}=0.15$ to enable the frame $x_{0}$ to be more involved in the spatial correlation control of the current frame for the sake of better appearance preservation. Apart from this, retaining a relatively small portion of feature interaction for the current frame and the preceding frame simultaneously is promised to enhance motion stability and improve temporal coherence of the generated animation.
Furthermore, it is vital to note that we do not replace all the U-Net Ronneberger et al. (2015) transformer blocks with DCAM. We find that incorporating the DC Attention only in the upsampling blocks of the U-Net architecture while leaving the remaining unchanged allows us to maintain consistency with the identity and background details of the source, without compromising the current frame’s pose and layout.
@@ 3.4 Mask-Guided Decoupling Module
Directly utilizing the entire image features for attention fusion can result in a substantial loss of fine-grained details. To address this problem, we propose the mask-guided decoupling module, which decouples the character and background and enables individual inter-frame interaction to further refine spatial feature perception.
For the source image $I_{s}$, we obtain a precise body mask $M_{s}$ (i.e. $M_{x_{0}}$) that separates the character from the background by an off-the-shelf segmentation model Liu et al. (2023a). The target pose prior is insufficient to derive body mask for each generated frame of the character. Considering the strong semantic alignment capability of cross attention layers mentioned in Prompt-to-prompt Hertz et al. (2022), we extract the corresponding body mask $M_{x_{i}}$ for each frame from the cross attention maps. With $M_{s}$ and $M_{x_{i}}$, only attentions of character and background within corresponding region are calculated, according to the mask-guided decoupling module as follows:
$\text{K}^{c}_{j}=M_{x_{j}}$ $\odot\text{K}_{j},\text{K}^{b}_{j}=(1-M_{x_{j}})\odot\text{K}_{j},$ (9)
$\text{V}^{c}_{j}=M_{x_{j}}$ $\odot\text{V}_{j},\text{V}^{b}_{j}=(1-M_{x_{j}})\odot\text{V}_{j},$
@@
$\text{DCA}^{c}_{i}$ $=\lambda_{1}*\text{CFA}^{c}_{i,0}+\lambda_{2}*\text{CFA}^{c}_{i,i-1}+\lambda_{3}*\text{CFA}^{c}_{i,i},$ (10)
$\text{DCA}^{b}_{i}$ $=\lambda_{1}*\text{CFA}^{b}_{i,0}+\lambda_{2}*\text{CFA}^{b}_{i,i-1}+\lambda_{3}*\text{CFA}^{b}_{i,i},$
@@ 4.1 Experiment Settings
We implement PoseAnimate based on the public pre-trained weights of ControlNet Zhang et al. (2023) and Stable Diffusion Rombach et al. (2022) v1.5. For each generated character animation, we generate $N=16$ frames with a unified $512\times 512$ resolution. In the experiment, we use DDIM sampler Song et al. (2020) with the default hyperparameters: number of diffusion steps $T=50$ and guidance scale $w=7.5$. For the pose-aware control module, loss function of optimizing text embedding $\varnothing_{text}$ is MSE. The optimization iterations are $250$ in total with $n=5$ inner iterations per step, and the optimizer is Adam. All experiments are performed on a single NVIDIA A100 GPU.
@@ 4.2 Comparison Result
We compare our PoseAnimate with several state-of-the-art methods for character animation: MagicAnimate Xu et al. (2023) and Disco Wang et al. (2023a). For MagicAnimate (MA), both DensePose Güler et al. (2018) and OpenPose signals of the same motion are applied to evaluate the performances. We leverage the official open source code of Disco to test its effectiveness. Additionally, we construct a competitive character animation baseline by IP-Adapter Ye et al. (2023) with ControlNet Zhang et al. (2023) and spatio-temporal attention Wu et al. (2023), which is termed as IP+CtrlN. It is worth noting that these methods are all training-based, while ours does not require training.
@@
We set up two different levels of pose for the experiments to fully demonstrate the superiority of our method. The visual comparison results are shown in Fig. 4, with the left side displaying simple actions and the right side complex actions. Although IP+CtrlN has good performance on identity preservation, it fails to maintain details and inter-frame consistency. Disco completely loses character appearance, and severe frame jitter leads to ghosting shadows and visual collapse for complex actions. MagicAnimate performs better than the other two methods, but it still encounters inconsistencies in character appearance at a more fine-grinded level guided by DensePose. It is also unable to preserve background and character details accurately, e.g., vehicle textures and masks of the firefighter and the boy in Fig. 4. MagicAnimate under OpenPose signal conditions has worse performances than that under DensePose. Our method exhibits the best performance on image fidelity to the source image and effectively preserves complex fine-grained appearance details and temporal consistency.
@@
Table 1: Quantitative comparison between PoseAnimate and other training-based state-of-the-art methods. The best average performance is in bold. ↑ indicates higher metric value and represents better performance and vice versa. MA stands for MagicAnimate.
For quantitative analysis, we first randomly sample 50 in-the-wild image-text pairs and 10 different disered pose sequences to conduct evaluations. In this section, we adopt four evaluation metrics: (1) LPIPS Zhang et al. (2018) measures the fidelity between generated frames and the source image. (2) CLIP-I Ye et al. (2023) represents the similarity of CLIP Radford et al. (2021) image embedding between generated frames and the source image. (3) Frame Consistency (FC) Esser et al. (2023) evaluates video continuity by computing the average CLIP cosine similarity of two consecutive frames. (4) Warping Error (WE) Liu et al. (2023b) evaluates the temporal consistency of the generated animation through the Optical Flow algorithm Teed and Deng (2020). Quantitative results are provided in Tab. 1. Our method achieves the best scores on LPIPS and CLIP-I, and significantly surpasses other comparison methods in terms of fidelity to the source image, demonstrating outstanding detail preservation capability. In addition, PoseAnimate outperforms two training-based methods in terms of interframe consistency and obtains a good Warping Error score, illustrating that it is able to ensure good temporal coherence without additional training.
To further make a comprehensive quantitative performance comparison, we also follow the experimental settings in MagicAnimate, and evaluate both image fidelity and video quality on two benchmark datasets, namely TikTok Jafarian and Park (2021) and TED-talks Siarohin et al. (2021). we compare FID-VID Balaji et al. (2019) and FVD Unterthiner et al. (2018) metrics for video quality, as well as two essential image fidelity metrics, L1 and FID Heusel et al. (2017). The experimental results are presented in Tab. 2, where PoseAnimate achieves state-of-the-art image fidelity while maintaining competitive video quality.
Method Image Video
[…]
IP+CtrlN 7.13E-04 68.23 93.56 724.37
DisCo 3.78E-04 30.75 59.90 292.80
MA 3.13E-04 32.09 21.75 179.07
Ours 3.06E-04 31.47 63.26 286.33
(a) Quantitative comparisons on TikTok dataset.
Method Image Video
[…]
IP+CtrlN 4.06E-04 45.75 38.48 281.42
DisCo 2.07E-04 27.51 19.02 195.00
MA 2.92E-04 22.78 19.00 131.51
Ours 1.98E-04 21.24 20.15 168.02
(b) Quantitative comparisons on TED-talks dataset.
Table 2: Quantitative performance comparison, with best performance in bold and second best underlined. MA corresponds to MagicAnimate (DensePose).
@@ 4.3 Ablation Study
Figure 5: Visualization of ablation studies, with errors highlighted in red circles. Source prompt: “An iron man on the road.”
We conduct an ablation study to verify the effectiveness of each component of our framework and present the visualization results in Fig. 5. The leftmost one in the first row is the source image and the others are the target pose sequences. The following rows are generation results without certain components: (a) Pose-Aware Control Module that effectively removes the interference of the source pose and maintains consistency of the content unrelated to the character; (b) Dual Consistency Attention Module that restores and preserves character identity while also improves temporal consistency; (c) Masked-Guided Decoupling Module that preserves fine-grained details and enhances animation fidelity; and (d) Pose Alignment Transition Algorithm that tackles the issue of pose misalignments while enabling smooth motion transitions.
@@
Fig. 5(a) illustrates the significant interference of the original pose on the generated actions. Due to the substantial difference between the posture of Iron Man’s legs in the source and in the target, there is a severe breakdown in the leg area of the generated frame, undermining the generation of a reasonable target action. Moreover, character-independent scenes also have noticeable distortion.
@@
From Fig. 5(b) we can find that it fails to maintain character identity consistency without DCAM. And the missing pole and Iron Man’s hand in the red circles reveal inter-frame inconsistency, indicating that both spatial and temporal consistency cannot be effectively maintained.
@@
Compared with our results in Fig. 5(e), it can be observed that small signs are missing without MGDM. It shows that MGDM can effectively enhance the perception of fine-grained features and image fidelity.
@@
Fig. 5(d) verifies the proposed Pose Alignment Transition Algorithm. The red circles in the second frame indicate the spatial content misalignment. When Iron Man in the original image does not match with the input pose position, an extra tree appears in the original position of Iron Man. And such misalignment can also lead to disappearance of background details, e.g., streetlights and distant signage.
@@ 5 Conclusion
This paper proposes a novel zero-shot approach PoseAnimate to tackle the task of character animation for the first time. Through the integration of three key modules and an alignment transition algorithm, PoseAnimate can efficiently generate high-fidelity, pose-controllable and temproally coherent animations for a single image across diverse pose sequences. Extensive experiments demonstrate that PoseAnimate outperforms the state-of-the-art training based methods in terms of character consistency and detail fidelity.
Acknowledgments
This project was supported by NSFC under Grant No. 62032006.